Dear Richard dear all,

please find below a short summary of the activities done today on the mountain related to the AO system. Italian people working at LBTO where:

Arcetri: E. Pinna, A. Puglisi, L. Fini, M. Sozzi, P. Stefanini
ADS: E. Anaclerio

W unit
1) WFS box transported to LBTO. The WFS board has been unpacked and visually inspected. No damage was found. The unit has been prepared for installation into the AGW unit.
2) the W unit HW mounted in the AGw unit has been unpacked and visually inspected. No damage was found.
3) The AGw/W unit has been powered from the main plug all together.
4) A functionality check of the following devices has been performed succesfully:
a) auxiliary unit bench stages and lamp
b) WFS ccd (CCD39) and acquisition camera CCD (CCD47) control electronics
c) BCUs unit
d) WFS positioning XYZ stages (Bayside)
e) temperature sensors and humidity sensors

AO control SW
1) Network VLAN was fixed in the morning. Network tests successful with copper wiring.
2) Communication with WFS internal devices tested locally in the lab with copper wiring.
3) Communication with F.O. from wfsdx server in control room also tested.
4) Completed the software configuration of the WFS part by adding the proper configuration and calibration files. A few configuration details modified to adapt to the local environment.

LBT672a
Unpacking of the LBT672 HW is started. E Anaclerio has been working to re-assemble the unit.

Next update tomorrow.

All the best

Simone

-- ArmandoRiccardi - 02 Apr 2010
Topic revision: r1 - 02 Apr 2010, ArmandoRiccardi
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback